OpenStudy (anonymous):

During a rain storm, the level of a river rose 1 1/4 inches in 4 1/2 hours. How many inches did it rise in one hour?

OpenStudy (anonymous):

@Directrix

OpenStudy (texaschic101):

(1 1/4) / (4 1/2) = x / 1 --- 1 1/4 inches in 4 1/2 hrs = x inches in 1 hr change mixed numbers to improper fractions (5/4) / (9/2) = x/1 cross multiply (9/2)x = 5/4 -- divide both sides by 9/2 x = (5/4) / (9/2) x = 5/4 * 2/9 x = 10/36 which reduces to 5/18 so it rose 5/18 of an inch in 1 hr

OpenStudy (anonymous):

@rishavraj